# The Serial Scaling Hypothesis

### v1

@@ Abstract
While machine learning has advanced through massive parallelization, we identify a critical blind spot: some problems are fundamentally sequential. These "inherently serial" problems—from mathematical reasoning to physical simulations to sequential decision-making—require dependent computational steps that cannot be parallelized. Drawing from complexity theory, we formalize this distinction and demonstrate that current parallel-centric architectures face fundamental limitations on such tasks. We argue that recognizing the serial nature of computation holds profound implications on machine learning, model design, hardware development. As AI tackles increasingly complex reasoning, deliberately scaling serial computation—not just parallel computation—is essential for continued progress.
@@ 1 Introduction
Scaling up machine learning has driven remarkable progress [1, 44, 26, 49, 55], but not all scaling is created equal. Although the field has focused on increasing parallel computation [23, 17, 87, 102, 46]—with bigger GPUs, wider models, and more data—some problems stubbornly resist these advances [2, 51, 59, 71, 69, 16]. For a class of tasks, only increasing serial computation—allowing models to perform more sequential steps—yields further progress.
@@
The recognition that some problems fundamentally require deep or sequential models has remained mostly a theoretical curiosity [115, 72, 13]. Only recently has there been a growing recognition of scaling test-time computation, in contrast to train-time computation [105, 82, 128]. Yet this dichotomy still overlooks the role of parallel vs. serial computation which is central to this paper. In practice, both parameter counts and computation FLOPs are largely treated as single numbers, with no distinction between width (parallel) and depth (serial computation) [49, 44, 82, 105].
@@
Consider Sudoku—a number-placement puzzle requiring each number 1–9 to appear once per row, column, and subgrid—as a parable (Figure 1). Easy puzzles can be solved by filling in many blanks independently, in parallel. Hard puzzles, however, require a long chain of dependent reasoning: each blank depends on the others. No algorithm can shortcut the process. This distinction—between problems that are “wide” (parallel) and those that are “deep” (inherently serial)—is fundamental; it dictates how we design efficient models, yet it is underappreciated in machine learning.
@@
We propose the Serial Scaling Hypothesis: For many important ML problems, especially those involving complex reasoning, planning, or the evolution of interacting systems, increasing parallel computation alone is insufficient. Progress requires scaling the amount of serial computation.
This hypothesis is motivated by both theory and practice. Complexity theory shows that some problems cannot be efficiently parallelized. Empirically, tasks such as hard Sudoku, mathematical reasoning, and sequential decision-making benefit disproportionately from models that can perform more serial steps (e.g., deeper networks, longer chains of thought, more search iterations).
@@
Why does this matter for machine learning? As we push toward more challenging tasks—advanced reasoning, physical simulations, planning, and scientific discovery—we encounter problems that parallel architectures (like Transformers and diffusion models) cannot efficiently solve. Recognizing the limits of parallel scaling and the necessity of serial computation has several implications:
@@
  • Model design: Should we revisit architectures that allow for deeper or more sequential computation?
@@
  • Evaluation: Should we measure and report serial compute separately from total compute?
Contributions. We (i) formulate the Serial Scaling Hypothesis and connect it to complexity theory, (ii) survey a range of problems—cellular automata, many-body physics, sequential decision-making, and advanced reasoning—where serial computation is provably or empirically essential, (iii) analyze the limitations of modern parallel-centric machine learning models on these tasks, and (iv) discuss implications for future model and hardware design, with practical recommendations.
1.1 A Parable: Easy vs. Hard Sudoku
Figure 1: The four circled blanks in the easy Sudoku have no interdependence and can be filled in parallel. The blanks in a hard Sudoku puzzle are densely interdependent, such that one blank cannot be filled conclusively until one has worked out a long chain of consequences for many others.
Not all hard problems are hard in the same way. An easy Sudoku puzzle can be solved by filling in many blanks independently—each can be completed in parallel. In contrast, a hard Sudoku requires a long chain of dependent reasoning: each blank depends on the others, and the solution must be built step by step. No algorithm exists that can shortcut this process; only more sequential computation—more "depth"—helps. While a large number of easy puzzles may take the same amount of computation as a single hard puzzle, yet one can be significantly accelerated by more processors, while the other cannot. This distinction between “wide” (parallelizable) and “deep” (inherently serial) computation is central to our thesis. Many tasks in machine learning, from advanced reasoning to planning, share this serial character: they require a sequence of dependent steps, not just more parallel computation.
1.2 A Brief History of Scaling

### v2

@@ Abstract
While machine learning has advanced through massive parallelization, we identify a critical blind spot: some problems are fundamentally sequential. These "inherently serial" problems—from mathematical reasoning to physical simulations to sequential decision-making—require sequentially dependent computational steps that cannot be efficiently parallelized. We formalize this distinction in complexity theory, and demonstrate that current parallel-centric architectures face fundamental limitations on such tasks. Then, we show for first time that diffusion models despite their sequential nature are incapable of solving inherently serial problems. We argue that recognizing the serial nature of computation holds profound implications on machine learning, model design, and hardware development.
@@ 1 Introduction
The scaling up of machine learning has driven remarkable progress (Achiam et al., 2023; Hoffmann et al., 2022; Dosovitskiy et al., 2021; Kaplan et al., 2020; Krizhevsky et al., 2012), much of this has come from parallel scaling: hardware shifted from CPUs to massively parallel GPUs, architectures moved from RNNs to highly parallelizable Transformers, and algorithms increasingly exploit parallel compute (Dao, 2024; Chowdhery et al., 2023; Jouppi et al., 2017). But some problems stubbornly resist such advances (Aggarwal & Welleck, 2025; Kazemi et al., 2025; Li et al., 2024; Merrill & Sabharwal, 2023b; Marcus, 2018; Chollet, 2019): for a class of tasks, only scaling serial computation—allowing models to perform more sequential steps—yields further progress. It seems not all scaling is created equal.
@@
The necessity of deep or sequential models for some problems has remained mostly theoretical (Telgarsky, 2016; Merrill & Sabharwal, 2025; Chen et al., 2024a). Only recently has scaling test-time computation become recognized, independently of train-time computation (Snell et al., 2024; OpenAI, 2024; Wu et al., 2025). Yet this dichotomy still overlooks the role of parallel vs. serial computation which is central to this paper. The literature on scaling still commonly report both parameter counts and computation FLOPs each as a single number, treating width (parallel) and depth (serial computation) interchangeably (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2024; Snell et al., 2024).
Figure 1: (Left) Many easy Sudoku puzzles, where the circled blanks can be filled independently in parallel. (Right) A hard Sudoku with the same total compute, but the circled blanks are interdependent, requiring sequential reasoning.
@@
Consider Sudoku—a number-placement puzzle requiring each number 1–9 to appear once per row, column, and subgrid—as a parable (Figure˜1). Easy puzzles can be solved by filling in many blanks independently, in parallel. Hard puzzles, however, require a long chain of dependent reasoning: each blank depends on the others. No algorithm can shortcut the process.
Solving many easy puzzles may cost the same total computation as solving a single hard one, but only the easy ones can be sped up with more processors. This distinction—between problems that are “wide” (parallel) and those that are “deep” (inherently serial)—is fundamental, yet it is underappreciated in machine learning.
@@
The Serial Scaling Hypothesis (SSH) For many important ML problems such as reasoning, decision making, and modeling dynamic systems, increasing parallel computation alone is insufficient. Progress requires scaling the amount of serial computation.
The appeal of this hypothesis is that it is both theoretically sound and practically relevant:
  • Grounded in theory: Complexity theory proves some problems parallelize efficiently, while others do not (Greenlaw et al., 1995).
  • Explains past successes: The breakthrough of deep learning came from increasing network depth (LeCun et al., 2015; Prince, 2023), and Chain-of-Thought (CoT) improves performance by adding more serial steps (Kojima et al., 2022; Li et al., 2024; Merrill & Sabharwal, 2024).
  • Guides practice: The parallel–serial lens guides both theory and practice. Formal results show bounded serial capacity in Transformers (Merrill & Sabharwal, 2023b) and expanded capacity with CoT (Li et al., 2024; Merrill & Sabharwal, 2024), matching empirical gains (Aggarwal & Welleck, 2025; Muennighoff et al., 2025). Our new analysis extends this picture to diffusion: despite thousands of iterations, diffusion models with a $\mathsf{TC}^{0}$ backbone have only constant computation depth, consistent with observed step-count plateaus (Ma et al., 2025; Ravishankar et al., 2024).
@@
Why does this matter for machine learning? As we push toward more challenging tasks—advanced reasoning, physical simulations, planning, and scientific discovery—we encounter problems that parallel architectures (like Transformers and diffusion models) cannot efficiently solve. Recognizing this has several implications:
@@
  • Model: Should we revisit architectures that allow for deeper or more sequential computation?
@@
  • Task analysis: Can some inherently serial tasks be reformulated into other tasks with reduced serial structure, while remaining useful in practice?
Contributions. We (i) introduce the Serial Scaling Hypothesis (SSH), (ii) prove that diffusion models, despite many iterative steps, have limited serial capacity, (iii) prove that certain Markov decision problems require serial computation for good decisions, (iv) identify machine learning problems where serial computation is essential, and (v) discuss SSH implications for machine learning.
2 Machine Learning from the Serial Perspective
Figure 2: . (A) A decision problem has a variable-size input and a fixed-size output (e.g., “yes”/“no”). (B) A serial problem requires deeper or more steps as the problem size grows. Examples of serial problems are: (C) Cellular automaton: takes the initial state as input and outputs a discrete value of the row NN at cell $i$ for $i\in\{1,\dots,2N-1\}$. (D) Many-body mechanics: takes initial positions and momenta of each particle with time $T$ as inputs and outputs the particle locations at time $T$ in a limited-precision space. (E) Math QA: takes a question as input and outputs the answer autoregressively, with each output from a fixed set of possibilities.
We first formalize inherently serial problems (Section˜2.1), show they are not only theoretically valid but also pervasive in machine learning (Section˜2.2). We then examine why predominant ML models struggle with these tasks (Section˜2.3). Finally, we discuss broader implications for ML (Section˜2.4).
2.1 Complexity Framework for Inherently Serial Problems
We focus our attention on binary decision problems which take in $N$ input tokens and output either “yes” or “no” as depicted in Figure˜2(A). While seemingly limited, this framework includes all problems with discrete responses, since making a discrete choice out of $2^{n}$ options is equivalent to answering $n$ yes/no questions Figure˜2 (C–E) show how cellular automata, many-body mechanics, and open-ended question answering can be represented in this way.
We adopt the complexity class $\mathsf{TC}$ (Threshold Circuits) to formally distinguish between serial and parallel problems, since it is the standard theoretical framework that formalizes the serial–parallel divide, and it has strong connections to neural networks.
Definition 2.1 (Informal, see Appendix˜D).
A problem is in $\mathsf{TC}^{i}$ if, for inputs of size $N$, it can be solved by a Boolean circuit with polynomial width and polylogarithmic depth, using basic logic gates (AND, OR, NOT) plus majority gates. The class $\mathsf{TC}$ is the union of all such $\mathsf{TC}^{i}$ for $i\geq 0$.
Intuitively, a problem belongs to $\mathsf{TC}$ if and only if it can be solved by a multilayer perceptron (MLP) with polynomial width and polylog depth, where polylog denotes $\sf{poly}(\log N)$ (Parberry & Schnitger, 1988). A problem belongs to $\mathsf{TC}^{0}$ if and only if it can be solved by a constant-depth MLP. We consider $\mathsf{TC}$ problems as parallel problems, since there are sublinear-depth MLPs that solve them.
Figure 3: The complexity classes are nested as $\mathsf{TC}^{0}\subseteq\mathsf{TC}^{1}\subseteq\dots\subseteq\mathsf{TC}\subseteq\mathsf{P}$. Each containment is widely believed to be strict. Problems in $\mathsf{TC}$ are parallel, while those outside are inherently serial.
Although the universal approximation theorem (Cybenko, 1989; Hornik et al., 1989) shows that a 3-layer MLP with unbounded width can approximate any continuous function, a constant-depth MLP with only polynomial width is far more limited, capturing only problems in $\mathsf{TC}^{0}$.
A common phenomenon in computational complexity is that one can make a neural network shallower at the cost of exponentially larger width (Valiant, 1983; Hajnal et al., 1993; Sherstov, 2007; Williams, 2014; Oliveira & Santhanam, 2015; Eldan & Shamir, 2016; Liang & Srikant, 2016; Cohen et al., 2016; Telgarsky, 2016; Merrill & Sabharwal, 2025). However, such exponential-sized networks are intractable in both memory and computation. We consider them as inefficient solutions in this paper. This exponential depth–width trade-off underlines the importance of characterizing neural networks depth and width separately, which motivates the Serial Scaling Hypothesis.
Definition 2.2.
Assumption 2.3.
We adopt the widely held belief that $\mathsf{TC}\subsetneq\mathsf{P}$—that is, some polynomial-time problems are inherently serial (Greenlaw et al., 1995, Ch. 5, Ch. 8).
2.2 Real-world problems are likely inherently serial
A natural objection to theory: “Do such theoretical obstructions to parallelism arise in the real world?” We say: Yes—particularly if one works in domains involving the need for physical simulation or reinforcement learning. But even outside these areas, it remains highly likely that one will encounter inherently serial problems. Even apparently simple tasks with a polynomial-time algorithm can be inherent serial (Assumption 2.3).
Drawing from both empirical observations and theoretical arguments presented in the following Section˜3, we propose the following complexity-theoretic hypothesis:
Hypothesis 2.4 The following problems—cellular automata evolution, many-body mechanics, sequential decision problems, mathematical question answering—are inherently serial.
Inherently serial problems (see Definitions 2.1, 2.2) exhibit fundamental computational dependencies: the outcome of intermediate steps directly influences subsequent steps in ways that cannot be shortcut without compromising correctness. To make this intuition rigorous, we prove a new theorem (Theorem˜G.2) on inherent seriality in sequential decision making.
Limitations. (1) These arguments apply to the general (worst-case) complexity of the problems. A problem may be inherently serial in general, but a parallel algorithm may cover most specific instances occurring in practice. (2) An inherently serial problem may be solved approximately in parallel to an accuracy acceptable for practical use.
2.3 Are our models capable of solving inherently serial problems?
Method Parallel?
Solve serial
problem?
MLPs
Transformers
SSMs and Mamba
RNNs
Repeating layers
Chain-of-Thought
Diffusion models
($\sf{TC}^{0}$ backbone)
11footnotemark: 1
Most modern architectures are designed for scaling parallel computation. Transformers and state-space models (SSMs) (Gu et al., 2021; Gu & Dao, 2023), though often called “sequence” models, process all available input in parallel. This raises a crucial question: “How efficiently do they solve inherently serial problems?”
Prior work shows that for fixed-depth, polynomial width, and fixed-precision, various architectures, including MLPs, Transformers, and linear SSMs, all collapse into constant-depth threshold circuits ($\mathsf{TC}^{0}$) (Merrill et al., 2022; Chiang, 2024; Chen et al., 2024b; Merrill et al., 2024). Intuitively, their computational graphs all collapse into constant parallel steps, never requiring long sequential chains. Importantly, these results characterize the architectures themselves. A model, however, is the combination of both architecture and inference procedure. Thus, while a transformer architecture lies in $\mathsf{TC}^{0}$, a Transformer model run with serial inference—such as autoregressive CoT or recurrence—extends beyond $\mathsf{TC}^{0}$ and can capture inherently serial computations (see Section˜C.1). Table˜1 summarizes computational characteristics of common machine learning models.
Diffusion models cannot solve inherently serial problems. A key contribution of the present paper is a new theorem (Theorem˜F.1) on the limitations of diffusion modeling. Although the step-by-step sampling of a diffusion model appear serial, we prove in Section˜4 that a diffusion model with a $\mathsf{TC}^{0}$ backbone remains in $\mathsf{TC}^{0}$, even with infinitely many sampling steps. Thus, despite their stepwise structure, diffusion models are incapable of solving inherently serial problems.
By linking the computational limits of modern ML models with the inherently serial structure of many key problems, we arrive at the following conclusion:
Key Limitation of Modern Machine Learning Models MLPs, Transformers, SSMs, and diffusion models with $\mathsf{TC}^{0}$ backbones are provably incapable of solving general instances of inherently serial problems such as cellular automata evolution, many-body mechanics, sequential decision-making, and mathematical question answering.
Limitations. (1) Certain instances of serial problems may be trivial or admit parallel solutions; seriality holds only in the general case. (2) For problems beyond $\mathsf{P}$ (e.g., $\mathsf{NP}\setminus\mathsf{P}$), exponential computation—not serial computation—dominates. Thus, the Serial Scaling Hypothesis applies most directly to problems of practical real-world difficulty.
2.4 Implications of The Serial Scaling Hypothesis
For machine learning practitioners. Many important real-world problems—such as cellular automata, many-body mechanics, and sequential decision-making—are inherently serial. Without sufficient serial computation, solving them with shallow models requires an exponentially large set of weights, which in turn demands exponentially large datasets to train. Neither is affordable in practice. This mismatch—solving inherently serial problems using shallow or parallel models—helps explain why many ML systems generalize poorly beyond their training distributions (Torralba & Efros, 2011; Zhang et al., 2017; Recht et al., 2019; Liang et al., 2023; Mancoridis et al., 2025; Zhang et al., 2025).
Evidence from RL fine-tuning (Wang et al., 2025b; Yue et al., 2025; Shao et al., 2025) suggests that LLMs already acquire both shallow heuristics (fast but memorization-like (Nikankin et al., 2025)) and deeper algorithms (slower but more general) during pretraining. However, if the inference-time compute budget is set too low, the model defaults to the shallow routines, producing fast but brittle behavior. Only with sufficient serial computation, the deeper routines can be executed resulting in better generalization.
For task & benchmark designers. Recognizing the inevitable cost to inherent seriality, serial problems may be reformulated into coarser or approximate problems to reduce serial depth to acceptable levels while remain practically useful. For RL, truncated value functions cap the effective depth while retaining theoretical guarantees and practical utility (Park et al., 2025; Sutton & Barto, 2018; De Asis et al., 2019). For reasoning, complexity theory shows that coarse decision problems can be tractable even when exact solutions are not—for example, primality is in $\mathsf{P}$ while factoring remains hard (Agrawal et al., 2004).
In addition, benchmarks should have inherently serial problems as a separate category, to distinguish serial and parallel scaling in model performance.
For model designers. Solving these challenging real-world problems may require recurrent structures that increase serial computation alongside today’s predominantly parallel designs. However, recurrence and depth often amplify gradient variance (Bengio et al., 1994; Pascanu et al., 2013) and L-Lipschitzness (Bartlett et al., 2017; Fazlyab et al., 2019), making models harder to train. This motivates improved training techniques and novel architectures such as implicit gradients (Wang et al., 2025a), xLSTM (Beck et al., 2024), and test-time training (Sun et al., 2024).
For hardware designers. Massively parallel computing machinery, especially GPU clusters, enabled past progress in deep learning. Future progress in machine learning and computing in general [^2] Amdahl’s law (Amdahl, 1967; Gustafson, 1988) shows that seriality is a hard upper bound on speedups achievable by parallelism. also depends on progress in high-clockrate, sequential computing machinery, with reduced data movement overhead (Kang et al., 2021; Kaur et al., 2024).
2.5 Related Works
Our hypothesis is similar to the “parallelism tradeoff” hypothesis—that all parallelizable models, irrespective of design, must necessarily fail to solve inherently serial problems (Merrill & Sabharwal, 2023a; b). Our work builds on classical complexity theory, including the “depth-width tradeoff” (Vishkin & Wigderson, 1985), the “work” vs. “depth” contrast  Blelloch (1996), $\mathsf{P}$-completeness (Greenlaw et al., 1995, Ch. 8), and computational irreducibility (Wolfram, 2002).
While these ideas are well-established in complexity theory, our work brings them into the machine learning context—extending them to real-world tasks such as sequential decision-making and question answering, and highlighting the gap between these tasks and the limited serial capabilities of modern machine learning models.
3 Serial Problems
In this section, we highlight representative inherently serial problems: cellular automata (Section˜3.1), physical systems (Section˜3.2), and complexity-theoretic hardness results (Section˜3.3), as well as practical domains such as sequential decision making (Section˜3.4) and reasoning QA (Section˜3.5). These cases illustrate serial bottlenecks that readers may encounter in other domains.
3.1 Cellular Automata
Figure 4: A single run of Rule 110. Given the top row, the CA evolves row-by-row according to the 8 rules.
We begin with predicting the outcome of cellular automata (CA) (Sarkar, 2000; Codd, 2014), shown in Figure˜2(C), as a simple problem that is inherently serial. A CA is a grid of cells, each in one of finitely many states, updated in discrete steps according to a local rule based on its neighbors. Despite this simplicity, CA can exhibit behaviors ranging from predictable to complex dynamics such as Rule 110 (Figure˜4), which hints at inherent seriality of this problem.
Cellular automata are inherently serial. Rule 110 has been proven Turing-complete (Woods & Neary, 2009), so computing the state of a cell $x_{i}$ at row $N$ requires simulating each step in sequence, without shortcuts (Greenlaw et al., 1995, p. 58). This is not unique to Rule 110: many CA problems are $\sf{P}$-complete (Moore, 1997), i.e., not efficiently parallelizable.
What does this mean? Even systems governed by just 8 simple rules can forbid shortcuts: reducing serial computation often incurs an exponential parallel cost, showing that serial and parallel computation are not interchangeable.
3.2 Many-Body Mechanics
The previous section showed that inherent seriality can arise even in systems governed by a handful of local rules. We now turn to a more realistic setting: Newtonian many-body dynamics.
Consider $N$ particles evolving in $\mathbb{R}^{d}$ under forces and/or hard collisions. Given initial positions and momenta at $t=0$, the goal is to predict the particle positions at some later time $t=T$ in a finite-precision representation. [^3] Finite precision reflects practical scientific computation. With unbounded precision, related prediction problems become $\mathsf{PSPACE}$-hard. This is the instance depicted in Figure˜2(D).
Many-body mechanics is inherently serial. Classical mechanics is expressive enough to emulate arbitrary computation. The billiard-ball computer of Fredkin & Toffoli (1982) recreates any Turing machine using nothing but billiard balls elastically reflecting off each other and walls, and Moore (1990) proposes a similar recreation using smooth particle motion in a potential field in $\mathbb{R}^{3}$. Since the problem of simulating general Turing machines is inherently serial (Greenlaw et al., 1995, p. 58), an algorithm that accurately simulates general physical systems can only be done step by step.
Seriality in video prediction. Because videos capture inherently serial dynamics, such as collisions, forecasting the next frame from a sequence of $N$ frames is similarly inherently serial. This task underlies large-scale models for content creation (DeepMind, 2025; Brooks et al., 2024; Wan et al., 2025) and decision making (Bar et al., 2025; Yang et al., 2024).
The computational difficulty in video prediction does not come from the cost of rendering a single next frame. If a model maintains a complete world state (e.g., object positions and momenta), stepping that state forward is as efficient as applying a local physical law. The problem arises when losing track on the state: recomputing it requires replaying the chain of many-body mechanics from the last reliable observation, thus inherently sequential.
Figure 5: Predicting the frame at time $T$. The intermediate frames may not be observable by camera motion/occlusion.
In real videos, objects frequently leave the field of view or become occluded, as illustrated in Figure˜5. Current large-scale video predictors—typically Transformer-based and trained with next-frame objectives—are optimized only for immediate reconstruction. As a result, they are not encouraged to maintain the state of occluded objects. This can lead to an incomplete world state, effectively rendering the task inherently serial and no longer solvable in a single forward pass by the model. This may help explain why state-of-the-art video generators often produce physically inconsistent results (Kang et al., 2024; Brooks et al., 2024; Wan et al., 2025).
What does this mean? Tasks involving the modeling of system dynamics, including physical simulations and video prediction, likely require serial models to be efficiently solved. For video prediction, the model must maintain the full evolving world state.
3.3 $\mathsf{P}$-complete Problems
The inherently serial problems given previously have this in common: Given a problem statement of size $\mathcal{O}(n)$, it is immediately obvious how to construct a linear computational graph of length $\mathcal{O}(n)$, such that each node of the graph depends, and only depends, on the previous ones, and each step takes constant time. This matches the intuition of what being “serial” means.
Such problems have been formalized as $\mathsf{P}$-complete problems, of which the most prototypical one is the Circuit Value Problem (CVP) (Ladner, 1975): Given a Boolean circuit with a single output bit, specified as a graph of logic gates their connections, and an input binary string, what is the output bit? The CVP is robust in that many variations remain $\mathsf{P}$-complete (Greenlaw et al., 1995).
The obvious solution is to perform a topological sort (takes $\mathcal{O}(n)$ time) on the graph of the logic gates, so that each gate depends only on the output values of the previous gates, then evaluate them serially. The CVP embodies the intuition of what an inherently serial problem should be, which is that each step easily depends on the previous steps, but skipping steps is difficult. Since any $\mathsf{P}$ problem is efficiently reducible to the CVP, it is $\mathsf{P}$-complete.
This gives us an intuitive rule: If the problem appears to have a linear computational graph similar to the CVP, then it is likely inherently serial. Indeed, this intuition guided us in proving the theorem in Section˜3.4. Furthermore, an extensive list of problems are shown to be $\mathsf{P}$-complete (Greenlaw et al., 1995), thus suggesting that inherently serial problems are common in the wild.
3.4 Sequential Decision Problems
The goal in sequential decision problems is to obtain an optimal policy $\pi^{*}(s)=\arg\max_{\pi}J(\pi)$, where $J(\pi)$ is the expected return under policy $\pi$ in a Markov Decision Process (MDP) with finite horizon $N$, and discount factor $\gamma\in[0,1]$. Therefore, given a state $s$, the policy outputs an optimal action $a$ from a limited possibilities (discrete action) or with finite precision (continuous action). There are two aspects in this: finding a policy, and executing a policy.
Executing a policy is inherently serial. In Appendix˜G, we construct certain MDPs and prove that in these, any approximately optimal policy is inherently serial.
Finding a policy is inherently serial. Consider policy gradient methods (Sutton et al., 1999; Williams, 1992), including popular variants like PPO (Schulman et al., 2017), widely used in applications such as LLM fine-tuning (Ouyang et al., 2022).
In policy gradient, the parameters of a policy model are improved via gradient descent on return estimates, which must be unbiased for convergence to the optimal policy [^4] Convergence to suboptimal policies is still possible with biased return estimates (Mu & Klabjan, 2024; Tian et al., 2023).. Here, we show how an unbiased return estimation is inherently serial, implying that with parallel estimation the optimal policy is not guaranteed.
(a) MCTS in Hex board game: Performance improves with more MTCS expansion nodes across all training regimes. Perfect play is only possible with test-time MCTS. Data from Villalobos & Atkinson (2023).
(b) Actor & critic networks’ depth vs. width in locomotion & manipulation. A 2× deeper network (8 layers, width 256) outperforms a 16× wider network (4 layers, width 4,096) in Humanoid, a locomotion task with the largest observation & action spaces among the three. Data from Kevin et al. (2025).
If parallel return estimation is possible, one must be able to access state $i$ (and thus reward $r_{i}$) in fewer than a linear function of $i$ steps. This is unlikely in real-world MDPs, since even simple CA transition rules (Section˜3.1) or physical interactions (Section˜3.2) prevent such shortcuts [^5] As usual, there is an exponential tradeoff available. One can randomly generate a trajectory and check its consistency with the MDP and policy in parallel, but the expected number of trajectories needed before a consistent one is found grows exponentially with length $n$.. While scaling parallel computation via aggregating multiple trajectories can accelerate convergence by reducing variance (Sutton & Barto, 2018, p. 93), accurate return estimation still requires serial computation.
This inherent seriality motivates model-based reinforcement learning, where returns are computed by unrolling an internal model step by step. For instance, Monte Carlo Tree Search (MCTS)—which increases serial computation via tree expansion and reduces return-estimation bias—has achieved superhuman-level performance in diverse board games (Silver et al., 2016; 2017; Schrittwieser et al., 2020). Villalobos & Atkinson (2023) demonstrate consistent improvements in Hex from increasing MCTS expansion nodes (see Figure˜6(a)). Even in model-free RL, adding serial computation with deeper networks significantly outperform wider (more parallel) ones on locomotion tasks Kevin et al. (2025) (see Figure˜6(b)).
What does this mean? Parallel computation cannot substitute for serial computation in RL. Without sufficient serial computation, the optimal policy is not guaranteed.
3.5 Reasoning Question Answering
Figure 7: Average benchmark scores over 4 math benchmarks for longer reasoning chains (in different colors) vs. majority voting (as dots along each line). Data from Aggarwal & Welleck (2025).
Instead of going through an MDP as in RL, math question answering exhibits seriality through step-by-step logical reasoning, irrespective of the answer length, before arriving at the solution. As shown in Figure˜2(E), given a question as input tokens, the model autoregressively generates the solution, each step from a limited set of tokens.
Math QA is likely inherently serial. Solving grade-school mathematics, GSM8K (Cobbe et al., 2021), which has been used to benchmark reasoning capabilities in LLMs, can be formalized as dependency graphs (Ye et al., 2024). A solution is to traverse the graph in topological order and perform necessary arithmetic operations sequentially (Ye et al., 2024). This resembles the Arithmetic Circuit Value Problem (Greenlaw et al., 1995, p. 124), a generalization of the standard CVP to arithmetic operations. Like CVP, this problem is $\mathsf{P}$-complete—that is, inherently serial. Since seriality arises even in simple math QA, it likely extends to advanced benchmarks such as AIME and Olympiad-bench, where finding the correct approach is hard.
In mathematics QA, as shown in Figure˜7, Aggarwal & Welleck (2025) demonstrate that sequential scaling with longer reasoning chains consistently outperforms parallel scaling via majority voting controlled for the same token budget. This pattern holds across mathematical datasets of varying difficulty, including AMC, MATH, AIME, and Olympiad-bench.
Similarly, science QA also appears to exhibit inherent seriality. Muennighoff et al. (2025) report that in GPQA Diamond (Rein et al., 2024), a QA benchmark on PhD-level science, sequential scaling yields consistent accuracy improvements, only limited by the model’s context window, and is significantly more efficient than parallel scaling (via majority voting), which plateaus.
What does this mean? Complex question answering similar to math QA likely requires constructing answers step by step over a computation graph, and thus inherently serial.
4 Diffusion model’s computation is not serial
The Serial Scaling Hypothesis provide a lens to look at machine learning from serial–parallel lens. We have examined serial machine learning problems. We now turn to a diffusion models (Ho et al., 2020; Song et al., 2021; Song & Ermon, 2019), widely used in image/video generation (Brooks et al., 2024) and other vision tasks such as depth estimation (Saxena et al., 2023), and language modeling tasks (Li et al., 2022; Nie et al., 2025; Arriola et al., 2025). While commonly thought as a serial model, we show here for the first time that they are not.
Consider a problem with input $x_{1},\dots,x_{N}$ and fixed output $x_{N+1}$ (as in Figure˜2). A diffusion model uses a backbone neural network $\theta$ which takes the inputs and a noisy version of $x_{N+1}$, denoising it over $T$ steps until it becomes the output (see Figure˜8). This models the conditional distribution $p_{\text{truth}}(x_{N+1}\mid x_{1},\ldots,x_{N})$ or concisely $p_{\text{truth}}$.
While the number of denoising steps $T$ is scalable—more denoising steps for finer approximation of $p_{\text{truth}}$, in practice, diffusion models converge rapidly (Ma et al., 2025). For instance, image generation plateaus at 300 steps (Nichol & Dhariwal, 2021), depth estimation shows little difference between 5 and 100 steps (Ravishankar et al., 2024), and language modeling yields similar perplexity for 32 vs. 1024 steps (Austin et al., 2021). With distillation, 1–4 steps suffice without much loss (Yin et al., 2024; Liu et al., 2024; Song et al., 2023; Lin et al., 2024; Salimans & Ho, 2022). Such rapid convergence suggests that the effective computation depth of diffusion models is low. It would be surprising if the underlying computation were truly serial.
Diffusion models with a $\mathsf{TC}^{0}$ backbone can only solve problems in $\mathsf{TC}^{0}$. Previous work showed that a backward diffusion process converges to $p_{0}$ at the rate $TV=\mathcal{O}(d/T)$ (Li & Yan, 2024, Thm. 1), where $TV$ is the total variation between $p_{0}$ and the denoising $p_{\theta,0}$, and $d$ is the dimension of $x_{N+1}$. Building on top of it, we obtained the following theorem, the formal statement and proof of which are in Appendix F:
Theorem 4.1 (Informal).
What does this mean? Diffusion models only provide a constant amount of additional serial computation. The above theorem precludes the use of diffusion models as a scalable means of increasing serial computation. Unlike CoT, which genuinely adds serial compute (see Appendix˜E), diffusion models do not.
Limitation. The theorem doesn’t apply: (1) If the dimension of $x_{n+1}$ is increasing. However, diffusion language modeling, as output length increases, has yet to show promising results (Austin et al., 2021; Lou et al., 2024; Gulrajani & Hashimoto, 2023; Sahoo et al., 2024). (2) If the backbone is either poorly trained or trained under a different objective. This leaves open the possibility of serial-compute-scalable generative models beyond score-based diffusion.
5 Acknowledgements
We are grateful to William Merrill and Anant Sahai for their invaluable technical feedback and corrections that significantly strengthened our theoretical analysis. We also thank Angjoo Kanazawa, Justin Kerr, Amil Dravid, Yossi Gandelsman, and Yuatyong Chaichana for their thoughtful comments and suggestions that helped improve the clarity and presentation of this work. KP is supported by the ONR MURI grant.
6 Reproducibility
No new data or code has been used to support this paper. Some new theorems are introduced in this paper, and their proofs are in the appendix.
Appendix A Limitations
Our conclusions rely on the widely-held but unproven assumption that $\sf{TC}\subsetneq\sf{P}$ (˜2.3). If this assumption is disproven, our formalization of the serial/parallel dichotomy would be invalid. Moreover, our theoretical arguments apply only to the general case; although the intuition may extend to practical settings, such generalization is not guaranteed. In particular, the average-case complexity (Bogdanov et al., 2006) of many problems could differ or even allow parallel solutions.
For especially hard problems (e.g., $\mathsf{NP}\setminus\mathsf{P}$ and beyond), exponential cost—rather than serial depth—may dominate as the primary bottleneck. Thus, our focus is on real-world problems of practical difficulty.
Appendix B A Brief History of Scaling